OpenStudy (anonymous):

Calculate the present value of each of the following: a. $10000 in five years at 8% per annum compunded annually b. $5000 in three years at 9% pe annum compounded monthly

OpenStudy (anonymous):

Compound Interest. Make sure you remember this equation. Whatever helps to remember, use it. \[A=P(1+r)^n\] Where: "A" is the amount compounded "P" is the amount invested. "r" is the rate at which the money is compounded "n" is the period of time that the money was compounded

OpenStudy (anonymous):

a)\[A=$10000(1+0.08)^{5}\]

OpenStudy (anonymous):

but won't we use the geometric sequence for mula to solve?

OpenStudy (anonymous):

Okay do you have the answers to part a and b? If you do, can you show me the answer to part a please?

OpenStudy (anonymous):

a. 6805.83 b. 3820.74 i tried the formula that you suggested and it gives me the right answer but just curious to why the geometric sequence formula doesn't work in this case?

OpenStudy (anonymous):

oh! i used P=10000/(1.08)^5 and it gives me thr right answer

OpenStudy (anonymous):

Geometric formula?

OpenStudy (anonymous):

GP? ar^n-1?

OpenStudy (anonymous):

Sn=a(r^n-1)/(r-1) Sn= sum of n number of terms n= the number of terms r= common ratio its used for annuities

OpenStudy (anonymous):

Ah okay. the question befuddled me for a second. In this question, you're not finding the total amount/sum of the amount. You're finding the amount invested.

OpenStudy (anonymous):

The amount you originally put in to get $10000

OpenStudy (anonymous):

The sum of GP does work.

OpenStudy (anonymous):

\[S _{n}=\frac{ a(r^n-1) }{ r-1 }\] In that, you're trying to find a.

OpenStudy (anonymous):

Oops I wrote the wrong one. It's meant to be 1-r^n and 1-r

OpenStudy (anonymous):

because r<1, that's why you got the wrong answer.

OpenStudy (anonymous):

@hellomiss

OpenStudy (anonymous):

lemme try that one

OpenStudy (anonymous):

\[S _{n}=\frac{ a(1-r^n) }{1-r }\]

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no worries.